# Synergizing Foundation Models and Federated Learning: A Survey

### v1

@@ 4.1.3 Zeroth-Order Optimization
In contrast to the use of gradient descent in most FL optimization algorithms, a particular line of research advocates for the removal of BackPropagation (BP) Malladi et al. (2023a) in favor of Zeroth-Order Optimization (ZOO) Fang et al. (2022); Li and Chen (2021). BP-free methods conserve memory needed for computing gradients and minimize communication overhead for model aggregation Qin et al. (2024), making FMs more accessible for lower-end devices, thereby enhancing their applicability in diverse hardware environments.
@@
^$\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta};\mathcal{B})}{2\epsilon}\bm{z},$ (2)
where $\bm{z}\in\mathbb{R}^{d}$ with $\bm{z}\sim\mathcal{N}(0,\bm{I}_{d})$ and $\epsilon$ is the perturbation scale  Duchi et al. (2015). It requires only two forward passes through the model to compute the estimation of gradient, serving as a memory-efficient alternative to BP. However, Eq. (2) provides a biased gradient estimation, leading to a certain degree of information loss Liu et al. (2020). Alternatively, many studies opt for two-point gradient estimators that can yield a more stable and reliable approximation Spall (1992); Malladi et al. (2023a); Lin et al. (2023); Ling et al. (2024). The standard two-point gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
^$\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon\bm{z};\mathcal{B})}{2\epsilon}\bm{z}.$ (3)
Based on the above gradient estimation frameworks, recent work, such as that by Xu et al. (2024a); Lu et al. (2023b), has initiated preliminary explorations into the deployment of both FedPEFT and full-model fine-tuning of billion-sized FMs, like LLaMA, on mobile devices. The naive ZOO methods remain impractical for training large FMs in standard FL frameworks such as FedAvg, as they still result in a significant communication burden for model aggregation. In light of this, FedKSeed Qin et al. (2024) was proposed to further reduce communication overheads between the server and clients by using just a few random seeds and scalar gradients, requiring only a few thousand bytes for communication.
Although ZOO methods have shown promise in resource-efficient FL Ling et al. (2024), they generally require many iterations to achieve strong performance Malladi et al. (2023b). Compared to the well-established BP-based optimization, ZOO is still in the early stages of development, particularly for FM-FL settings, necessitating further research and optimization.
4.2 Adaptability
Adaptation refers to the process of tailoring a pre-trained FM to perform effectively across varying FL settings and scenarios. This mainly includes the capability to learn from different domains, cater to individual user needs, and work across diverse devices while retaining overall performance and efficiency. We focus on three key aspects of adaptation, namely domain-centric adaptation, client-centric adaptation, and system-centric adaptation.
4.2.1 Domain-Centric Adaptation
@@
Domain-Adaptive Pre-Training
@@
Despite being heavily reliant on large-scale and public datasets for their initial training, FMs often require further Domain-Adaptive Pre-Training (DAPT) with domain-specific data for tasks that necessitate specialized knowledge Gururangan et al. (2020); Guo and Yu (2022). In domains like healthcare, FL allows for the continued pre-training of these models using sensitive, domain-specific data without compromising privacy. Based on this idea, Jiang et al. (2023b) proposed FFDAPT, a computational-efficient further pre-training algorithm that freezes a portion of consecutive layers while optimizing the rest of the layers. Similarly, Wang et al. (2023) proposed FEDBFPT that builds a local model for each client, progressively training the shallower layers of local models while sampling deeper layers, and aggregating trained parameters on a server to create the final global model.
Multi-Domain Adaptation
@@
Given that client data may belong to various domains in real-world FL scenarios, some efforts Feng et al. (2023c); Su et al. (2024) have been devoted to facilitating multi-domain collaborative adaptation. Feng et al. (2023c) applied a pre-trained CLIP to the multi-domain scenario and proposed an adaptive prompt tuning method that uses domain-specific keys to generate prompts for each test sample. Furthermore, Su et al. (2024) employed knowledge distillation to selectively distill global knowledge based on an entropy measure, improving the generalization across different domains.
4.2.2 Client-Centric Adaptation
Client-centric adaptation refers to the process of tailoring an FM to meet the specific needs or preferences of individual clients while leveraging the decentralized and privacy-preserving nature of FL. Particularly, we discuss two types of popular personalized methods as follows:
Personalization
Adapter-based methods introduce small, trainable adapters into the frozen pre-trained FMs, allowing for client-specific model adaptation without altering the original FL. FedDAT Chen et al. (2024) leverages a dual-adapter structure, with personalized adapters focusing on client-specific knowledge and a global adapter maintaining client-agnostic knowledge. FedDAT executes bi-directional knowledge distillation between personalized adapters and the global adapter to regularize the client’s updates and prevent overfitting. Prompt-based methods involve using client-specific soft prompts to guide the model’s response. pFedPG Yang et al. (2023a) trains a prompt generator to exploit underlying client-specific characteristics and produce personalized prompts for each client, thereby enabling efficient and personalized adaptation.
Client Clustering
This branch of study aims to cluster clients based on the underlying relationships and tailor FMs for the client group with similar data distributions, thus reducing the negative impact of data heterogeneity and improving accuracy. Guo et al. (2024b) proposed a FedPEFT-based framework for multilingual modeling, which employs language family clustering to alleviate parameter conflicts of LoRA tuning.
4.2.3 System-Centric Adaptation
System-centric aims to improve adaptability at the system level. This involves handling resource heterogeneity in the FL systems while ensuring training efficiency and model utility.
Resource-Heterogeneous Methods
@@ 4.2.3 System-Centric Adaptation
Cross-device FL systems may be composed of devices equipped with heterogeneous resources, leading to disparities where certain devices exhibit more efficient model training than others Chen et al. (2024). To address this issue, several methods have been developed to customize model architectures for resource-heterogeneous FL systems. In FL environments possessing heterogeneous resources, LoRA-based FedPEFT exhibits distinctive flexibility and adaptation in fine-tuning frozen FMs without overburdening client devices. Su et al. (2023) suggested assigning LoRA adapters to varying numbers of layers for heterogeneous clients according to a randomly generated mask matrix. An alternative and more targeted idea is to choose diverse LoRA ranks across clients based on their system capabilities. Bai et al. (2024a) proposed FlexLoRA to adjust local LoRA ranks dynamically. FlexLoRA reconstructs the uniform full-sized LoRA module $\Delta\mathbf{W}$ for server-side model aggregation followed by an SVD-based parameter redistribution. However, concurrent research by Cho et al. (2024) has empirically demonstrated that the reconstruct-redistribute method suffers from performance loss compared to homogeneous LoRA. Instead, they proposed HetLoRA Cho et al. (2024) that utilizes zero-padding to align module size before aggregation. It then truncates the global LoRA modules for the specific rank of the next selected clients.
Split Learning
@@
Split learning addresses the resource heterogeneity between servers and clients by splitting a large model at a cut layer into client and server models Thapa et al. (2022). For each training step, the output tensor, so-called smashed data, from the client model and the corresponding labels are transmitted over to the server. The server continues the forward propagation by processing the smashed data through its remaining layers; it then computes the loss using the transmitted label and performs backpropagation. The gradient generated at the first layer of the server model is then transmitted back to the client for further backpropagation. Along this line, FedBERT Tian et al. (2022) proposes to leverage split learning for training the BERT model, showing the feasibility of training large FMs in FL settings. FedSplitX Shin et al. (2023b) is a more fine-grained method that allows multiple partition points for model splitting, accommodating more diverse client capabilities. Compared to conventional FL, split learning scales better with the size of FMs as it communicates only small-sized smashed data instead of model parameters Singh et al. (2019). Despite its merits, split learning is highly dependent on the network connection quality. Given that server-client interactions occur at every step of the optimization process Zheng et al. (2023), communication delays cause a more significant impact on efficiency.
4.3 Trustworthiness
This line of work aims to enhance trustworthiness throughout the FM-FL lifecycle, covering a variety of key aspects including, but not limited to, IP protection, privacy protection, and attack robustness.
4.3.1 IP Protection
@@
Existing IP protection involves safeguarding ownership of FMs from unauthorized use (e.g., model theft) Tekgul et al. (2021). We discuss the following two mainstream IP protection strategies: watermarking and black-box tuning.
Watermarking
@@
Watermarking is a well-known deterrence technology for model IP protection by providing the identities of model owners to demonstrate ownership of their models Adi et al. (2018). Tekgul et al. (2021) proposed WAFFLE, the first solution that addresses the ownership problem by injecting a watermark into the global model in FL environments. Recently, Yu et al. (2023b) proposed DUW that embeds a client-unique key into each client’s local model, aiming to identify the infringer of a leaked model while verifying the FL model’s ownership.
Black-Box Tuning
@@
Black-Box Tuning (BBT) is a set of ZOO-based methods that fine-tune FMs without direct access to model parameters Sun et al. (2022c, b). BBT methods are often additive, introducing additional parameters while keeping the original model frozen (see Section 4.1.1). Fed-BBPT Lin et al. (2023) is a general prompt tuning framework that facilitates the joint training of a global lightweight prompt generator across multiple clients. FedBPT Sun et al. (2024a) adopts a classic evolutionary-based ZOO method, CMA-ES Hansen and Ostermeier (2001), for training an optimal prompt that improves the performance of frozen FMs. ZooPFL Lu et al. (2023b), on the other hand, applies coordinate-wise gradient estimate to learn input surgery that incorporates client-specific embeddings. BBT allows for local fine-tuning of FMs while not infringing IP constraints. However, current research in this line is limited to few-shot learning with small datasets for LLM fine-tuning Sun et al. (2022b), while larger datasets and other modalities remain unexplored.
4.3.2 Privacy Protection
Protecting privacy in FM-FL requires both designing protective measures and studying privacy attack strategies.
Privacy-Preserving Techniques
Differential Privacy (DP) is a theoretical framework that governs privacy boundaries and manages the tradeoff between privacy and model convergence Wei et al. (2020); Xu et al. (2023). DP-based FL approaches often add artificial noise (e.g., Gaussian noise) to parameters at the clients’ side before aggregating to prevent information leakage Xu et al. (2023). Besides, DP is compatible with most FedPEFT methods. For instance, Sun et al. (2024b) showed that DP noise can even be amplified by the locally “semi-quadratic” nature of LoRA-based methods, motivating the integration of LoRA with DP to improve resource efficiency while maintaining data privacy Liu et al. (2023c). In addition to DP, Secure Multi-Party Computation (SMPC) Mugunthan et al. (2019) and Homomorphic Encryption (HE) Zhang et al. (2020) are also effective privacy-preserving mechanisms. However, they do not scale well enough for large-scale deployments in FM-FL.
Privacy Attack

### v2

@@ 4.3. Zeroth-Order Optimization
In contrast to the use of gradient descent in most FL optimization algorithms, a particular line of research advocates for the removal of BackPropagation (BP) (Malladi et al., 2023b) in favor of Zeroth-Order Optimization (ZOO) (Fang et al., 2022; Li and Chen, 2021). BP-free methods conserve memory needed for computing gradients and minimize communication overhead for model aggregation (Qin et al., 2024), making FMs more accessible for lower-end devices, thereby enhancing their applicability in diverse hardware environments.
Perturbation-based Methods.
@@
(2) $\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta};\mathcal{B})}{2\epsilon}\bm{z},$
where $\bm{z}\in\mathbb{R}^{d}$ with $\bm{z}\sim\mathcal{N}(0,\bm{I}_{d})$ and $\epsilon$ is the perturbation scale  (Duchi et al., 2015). It requires only two forward passes through the model to compute the estimation of gradient, serving as a memory-efficient alternative to BP. However, Eq. (2) provides a biased gradient estimation, leading to a certain degree of information loss (Liu et al., 2020). Alternatively, many studies opt for two-point gradient estimators that can yield a more stable and reliable approximation (Spall, 1992; Malladi et al., 2023b; Lin et al., 2023; Ling et al., 2024). The standard two-point gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
(3) $\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon\bm{z};\mathcal{B})}{2\epsilon}\bm{z}.$
Based on the above gradient estimation frameworks, recent work, such as that by Lu et al. (2023b); Xu et al. (2024a), has initiated preliminary explorations into both FedPEFT and full-parameter fine-tuning of billion-sized FMs, like LLaMA, on mobile devices. The naive ZOO methods remain impractical for training large FMs in standard FL frameworks such as FedAvg, as they still result in a significant communication burden for model aggregation. In light of this, FedKSeed (Qin et al., 2024) was proposed to further reduce communication overheads between the server and clients by using just a few random seeds and scalar gradients, requiring only a few thousand bytes for communication.
Others
Other resource-efficient techniques include model pruning and knowledge distillation. Model pruning reduces the size of a model by removing less important parameters, while knowledge distillation transfers knowledge from a larger model to a smaller one, enabling efficient deployment on resource-constrained devices.
Although ZOO methods have shown promise in resource-efficient FL (Ling et al., 2024), they generally require many iterations to achieve strong performance (Malladi et al., 2023a). Compared to the well-established BP-based optimization, ZOO is still in the early stages of development, particularly for FedFM settings, necessitating further research and development.
5. Adaptability
Adaptability refers to the capability of tailoring a pre-trained FM to perform downstream tasks across varying FL settings and scenarios. This mainly includes the ability to learn from different domains, cater to individual user needs, and work across diverse devices while retaining overall performance and efficiency. We focus on three key aspects of adaptation, namely domain-centric adaptation, client-centric adaptation, and system-centric adaptation.
5.1. Domain-Centric Adaptation
@@
5.1.1. Domain-Adaptive Pre-Training
@@
Despite being heavily reliant on large-scale and public datasets for their pre-training, FMs often require further Domain-Adaptive Pre-Training (DAPT) with domain-specific data for tasks that necessitate specialized knowledge (Gururangan et al., 2020; Guo and Yu, 2022). In domains like healthcare, FL allows for the continued pre-training of these models using sensitive, domain-specific data without compromising privacy. Based on this idea, Jiang et al. (2023b) proposed FFDAPT, a computational-efficient further pre-training algorithm that freezes a portion of consecutive layers while optimizing the rest of the layers. Similarly, Wang et al. (2023) proposed FEDBFPT that builds a local model for each client, progressively training the shallower layers of local models while sampling deeper layers, and aggregating trained parameters on a server to create the final global model.
5.1.2. Multi-Domain Adaptation
@@
Given that client data may belong to various domains in real-world FL scenarios, some efforts (Feng et al., 2024; Su et al., 2024b) have been devoted to facilitating multi-domain collaborative adaptation. Feng et al. (2024) applied a pre-trained CLIP to the multi-domain scenario and proposed an adaptive prompt tuning method that uses domain-specific keys to generate prompts for each test sample. Furthermore, Su et al. (2024b) employed knowledge distillation to selectively distill global knowledge based on an entropy measure, improving the generalization across different domains.
5.2. Client-Centric Adaptation
Client-centric adaptation aims to tailor FMs to meet the specific needs or preferences of individual clients while leveraging the decentralized and privacy-preserving nature of FL. This section covers three representative strategies for client-centric adaptation, namely personalization, client clustering, and preference-aware adaptation.
5.2.1. Personalization
Vanilla FL trains a single-globally shared model to fit the “average client”, which often struggles to capture the heterogeneous characteristics of individual clients. To mitigate this, personalization techniques prioritize adaptivity to the local client, tailoring the model to specific local distributions while preserving the benefits of collaborative learning.
Architecture-Based Personalization
By tailoring model architectures and the overall system designs, architecture-based personalization techniques decouple global knowledge sharing from local adaptation. Instead of training a monolithic model, these approaches often isolate specific parameter groups to exclusively capture user-specific patterns while aggregating the remaining parameters to maintain generalization. Leveraging FedPEFT, FedSA-LoRA (Guo et al., 2025) proposes a selective aggregation strategy where the $A$ matrix of the LoRA module learns global representations, while the $B$ matrix is retained locally to preserve personalization. Similarly, FedDPA (Yang et al., 2024b) introduces a dual-adapter architecture that explicitly maintains separate global and local adapters, dynamically fusing them via a gating mechanism during inference. Taking a generative approach to architectural design, pFedPG (Yang et al., 2023a) deploys a server-side prompt generator that dynamically produces client-specific visual prompts to adapt frozen FMs, thereby replacing direct parameter aggregation with personalized parameter generation. Instead of introducing additional learnable modules, FedSelect (Tamirisa et al., 2024) employs a dynamic masking mechanism to personalize the selection of sub-networks for local fine-tuning based on parameter importance. Furthermore, works like pFedLoRA (Yi et al., 2024) and FedCLIP (Lu et al., 2023a) leverage these modular designs to bridge model heterogeneity and align multi-modal features across diverse clients.
Optimization-Based Personalization
In contrast to structural modifications, optimization-based personalization focuses on tailoring the training dynamics and objective functions to accommodate local data distributions without necessarily altering the overall architecture. Instead of isolating specific parameters, these approaches guide the model toward personalized optimal points by adjusting the optimization trajectory, incorporating regularization terms, or modifying gradient updates. A prime example is FedRLHF (Fan et al., 2025a), which integrates Reinforcement Learning from Human Feedback (RLHF) into the federated workflow. By utilizing local reward models to evaluate generation quality, it steers the global policy to align with diverse user preferences through personalized reward signals. In terms of optimization stability, recent studies on ASR (Azam et al., 2023a) emphasize the critical role of optimizer-induced smoothness (e.g., using momentum-based optimizers) to navigate the non-IID loss landscapes of heterogeneous clients. Furthermore, approaches such as FedCLIP (Lu et al., 2023a) and communication-efficient strategies (Du et al., 2024) employ gradient correction mechanisms or multi-stage fine-tuning schedules to balance the trade-off between global convergence and local feature retention.
5.2.2. Client Clustering
Instead of tailoring models to individual clients, client clustering assigns specialized models to different groups of clients. The underlying assumption is that clients within a cluster possess high similarity in their data representations. A notable example is multilingual NLP, where clients exhibit linguistic heterogeneity characterized by inherent structural relationships, such as language families. Therefore, Guo et al. (2024c) proposed FedLFC, which explicitly groups clients by language family affiliations and assigns separate LoRA modules to each cluster to minimize parameter interference across distantly related languages. Extending this concept from flat clusters to a structured hierarchy, FedHLT (Guo et al., 2024b) organizes language adapters into a language tree structure for a coarse-to-fine parameter sharing scheme that captures both universal linguistic features and specific leaf-node characteristics. Complementary to these parameter-centric approaches, FedKC (Wang et al., 2022) suggests clustering within the latent representation space and exchanging “knowledge centroids” derived from local clusters, aligning the feature space across heterogeneous clients to handle skewed data distributions. Moving towards more dynamic grouping strategies, FL-TAC (Ping et al., 2024) proposes a task-specific adapter clustering mechanism; instead of grouping clients a priori, it clusters the learned low-rank adapters on the server side to identify and aggregate similar tasks data-drivenly. Furthermore, recent work (Chen et al., 2026) on sparsely-activated LLMs leverages a MoE architecture, which can be viewed as a form of dynamic parameter clustering, where diverse data patterns are automatically routed to specialized expert modules, achieving fine-grained adaptation without predefined partitions.
5.2.3. Preference-Aware Adaptation
In applications such as personalized chatbots, clients often desire tailored conversation styles and content that align with their individual preferences and needs. To achieve this, preference-aware adaptation techniques leverage user feedback, such as ratings or explicit preferences, to guide the fine-tuning process. Fan et al. (2025a) proposes FedRLHF, which integrates RLHF principles with federated reinforcement learning to address privacy and personalization challenges simultaneously. Similarly, Wu et al. (2025a) propose a preference-aware federated learning framework that incorporates user feedback to enhance the personalization of FMs. This approach allows the model to learn from user interactions and adapt its responses accordingly, leading to improved user satisfaction and engagement. Another example is the work by Wu et al. (2025a), which focuses on preference-aware fine-tuning of FMs using user feedback to optimize the model’s performance in generating personalized content. Additionally, Spadea and Seneviratne (2025) evaluate two preference-aware fine-tuning methods, i.e., Direct Preference Optimization (DPO) and Kahneman-Tversky Optimization (KTO), for LLMs in a federated setting. Their results indicate that KTO consistently outperforms DPO across various benchmarks, highlighting its flexibility and robust performance. They introduce a method that combines preference-aware fine-tuning with reinforcement learning techniques to further enhance the model’s ability to generate contextually relevant and personalized responses. This approach leverages user feedback as a reward signal in a reinforcement learning framework, allowing the model to learn from both explicit preferences and implicit feedback signals.
5.3. System-Centric Adaptation
System-centric adaptation focuses on enhancing the adaptability of FL systems from an architectural perspective. It primarily addresses the challenges arising from resource heterogeneity, communication efficiency, and system scalability when fine-tuning FMs in federated environments. As FL increasingly scales to cross-device and cross-silo deployments involving devices with diverse computational and communication capabilities, it becomes essential to design mechanisms that ensure efficient model training and aggregation under such constraints. This section reviews representative strategies for system-centric adaptation, including resource-heterogeneous methods and split learning frameworks, which jointly aim to maintain system-level efficiency and fairness without compromising model performance.
5.3.1. Resource-Heterogeneous Methods
@@
Cross-device FL systems may be composed of devices equipped with heterogeneous resources, leading to disparities where certain devices exhibit more efficient model training than others (Chen et al., 2024). To address this issue, several methods have been developed to customize model architectures for resource-heterogeneous FL systems. In FL environments possessing heterogeneous resources, LoRA-based FedPEFT exhibits distinctive flexibility and adaptation in fine-tuning frozen FMs without overburdening client devices. Su et al. (2024a) suggested assigning LoRA adapters to varying numbers of layers for heterogeneous clients according to a randomly generated mask matrix. An alternative and more targeted idea is to choose diverse LoRA ranks across clients based on their system capabilities. Bai et al. (2024) proposed FlexLoRA to adjust local LoRA ranks dynamically. FlexLoRA reconstructs the uniform full-sized LoRA module $\Delta\mathbf{W}$ for server-side model aggregation followed by an SVD-based parameter redistribution. However, concurrent research by Cho et al. (2024) has empirically demonstrated that the reconstruct-redistribute method suffers from performance loss compared to homogeneous LoRA. Instead, they proposed HetLoRA that utilizes zero-padding to align module size before aggregation. It then truncates the global LoRA modules for the specific rank of the next selected clients.
Given the success of MoE in scaling FMs efficiently (Jiang et al., 2024; Cai et al., 2025), recent research has pivoted towards adapting MoE for distributed environments (Wu et al., 2024). These efforts largely target inherent FL challenges such as data heterogeneity and resource heterogeneity. For instance, to capture diverse task patterns across clients, DGMoE (Miao et al., 2025) allows self-aware experts to adaptively determine their activation. To mitigate resource constraints, A3SMoE (Tran et al., 2025) and FFT-MoE (Hu et al., 2025) enable dynamic expert activation based on available client hardware. Furthermore, FedVLA (Miao et al., 2025) introduces an expert-driven aggregation strategy to ensure effective cross-client knowledge transfer. Recent works have adapted MoE for FL to address heterogeneity (Wu et al., 2024), such as DGMoE (Miao et al., 2025) for self-aware expert activation, and A3SMoE (Tran et al., 2025) or FFT-MoE (Hu et al., 2025) for hardware-dynamic activation.
5.3.2. Split Learning
@@
Split learning addresses the resource heterogeneity between servers and clients by splitting a large model at a cut layer into client and server models (Thapa et al., 2022). For each training step, the output tensor, so-called smashed data, from the client model and the corresponding labels are transmitted over to the server. The server continues the forward propagation by processing the smashed data through its remaining layers; it then computes the loss using the transmitted label and performs backpropagation. The gradient generated at the first layer of the server model is then transmitted back to the client for further backpropagation. Along this line, FedBERT (Tian et al., 2022) proposes to leverage split learning for training the BERT model, showing the feasibility of training large FMs in FL settings. FedSplitX (Shin et al., 2023b) is a more fine-grained method that allows multiple partition points for model splitting, accommodating more diverse client capabilities. Compared to conventional FL, split learning scales better with the size of FMs as it communicates only small-sized smashed data instead of model parameters (Singh et al., 2019). Despite its merits, split learning is highly dependent on the network connection quality. Given that server-client interactions occur at every step of the optimization process (Zheng et al., 2023), communication delays cause a more significant impact on efficiency.
6. Trustworthiness
In recent years, AI trustworthiness has been gaining attention from the government and different scientific communities, given the pressing need of ensuring AI systems operate reliably, safely, and in alignment with human values. For instance, the EU has proposed the AI Act (European Commission, 2021) to regulate AI systems based on their risk levels throughout their lifecycle. The distributed nature of FL introduces additional trustworthiness challenges and opportunities for FMs. In this section, we discuss three major aspects of trustworthiness in FedFM, including IP protection, privacy protection, and Byzantine Robustness.
6.1. IP Protection
@@
Existing IP protection involves safeguarding ownership of FMs from unauthorized use (e.g., model theft) (Tekgul et al., 2021). We discuss the following two mainstream IP protection strategies: watermarking and black-box tuning.
6.1.1. Watermarking
@@
Watermarking is a well-known deterrence technology for model IP protection by providing the identities of model owners to demonstrate ownership of their models (Adi et al., 2018). Tekgul et al. (2021) proposed WAFFLE, the first solution that addresses the ownership problem by injecting a watermark into the global model in FL environments. Recently, Yu et al. (2023b) proposed DUW that embeds a client-unique key into each client’s local model, aiming to identify the infringer of a leaked model while verifying the FL model’s ownership.
6.1.2. Black-Box Tuning
@@
Many FMs, particularly those designed for commercial use (e.g., ChatGPT and Gemini), often grant only black-box access by invoking APIs from service providers (Sun et al., 2022c). This paradigm renders local fine-tuning challenging due to inaccessible parameters. Recent research also explores federated black-box tuning methods to collaboratively customize FMs while maintaining the model parameters as inaccessible. For instance, a gradient-free prompt tuning framework, Fed-BBPT (Lin et al., 2023), has been proposed to free the clients from being required to access the model parameters. FedBPT (Sun et al., 2024a) ZooPFL (Lu et al., 2023b). Black-Box Tuning (BBT) is a set of ZOO-based methods that fine-tune FMs without direct access to model parameters (Sun et al., 2022c, b). BBT methods are often additive, introducing additional parameters while keeping the original model frozen (see Section 4.1.2). Fed-BBPT (Lin et al., 2023) is a general prompt tuning framework that facilitates the joint training of a global lightweight prompt generator across multiple clients. FedBPT (Sun et al., 2024a) adopts a classic evolutionary-based ZOO method, CMA-ES (Hansen and Ostermeier, 2001), for training an optimal prompt that improves the performance of frozen FMs. ZooPFL (Lu et al., 2023b), on the other hand, applies coordinate-wise gradient estimate to learn input surgery that incorporates client-specific embeddings. BBT allows for local fine-tuning of FMs while not infringing IP constraints. However, current research in this line is limited to few-shot learning with small datasets for LLM fine-tuning (Sun et al., 2022b), while larger datasets and other modalities remain unexplored.
6.2. Privacy Protection
Despite the privacy-preserving nature of FL, studies have shown that the distributed training process may still expose sensitive information to adversaries (Zhu et al., 2019), especially when the server is honest-but-curious (Boenisch et al., 2023), attempting to infer private data from shared model updates. This risk is exacerbated in the context of FMs due to their design to assimilate and generate contents from extensive, varied datasets, which inherently carry the risk of training data leakage (Yao et al., 2024a).
6.2.1. Privacy Attacks